OpenStudy (anonymous):

I am kind of confused about what I have to do to solve (3^4 mod 17)^2 mod 11 can anyone help me?

OpenStudy (mathmath333):

(3^4 mod 17)^2 mod 11 =(81 mod 17)^2 mod 11 =13^2 mod 11 =2^2 mod 11 =4

OpenStudy (mathmath333):

remember x mod y=a , here means a is the remainder when x is divided by y